Circular No. 8500

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN SAGITTARIUS S. Nakano, Sumoto, Japan, reports the independent discoveries of a

possible nova by Hideo Nishimura (Kakegawa, Shizuoka-ken, Japan; Pentax 6×7 camera + 200-mm f/4.0 lens + Kodak T-Max film, two exposures; new object at mag 8.7 on Mar. 28.779 UT) and by Yukio Sakurai (Mito, Ibakakiken, Japan; two CCD frames taken with a Fuji ‘Fine Pix S2’ digital camera + Nikon 180-mm f/2.8 lens; new object at mag 9.1 on Mar. 28.796). Nakano provides his measurement for the position of the new object from Sakurai’s jpeg image: α = 18^{h}16^{m}59^{s}.04, δ = −25^{o}56′38′′.8 (equinox 2000.0). Nothing was visible at this location on Sakurai’s CCD frames taken on Mar. 19, 20, and 25 (limiting mag 11.8). H. Yamaoka, Kyushu University, notes that nothing is visible at this location on an ASAS-3 image taken on Mar. 27.464 (limiting mag ∼ 14), adding that there is a bright infrared source in the 2MASS catalogue with position end figures 58^{s}.67, 39′′.0.

SUPERNOVA 2005ay IN NGC 3938 D. Rich, Hampden, ME, reports his discovery of an apparent supernova

(mag ∼ 15.6) on unfiltered CCD frames taken with a 0.31-m reflector on Mar. 27.08 and 27.99 UT. The new object is located at α = 11^{h}52^{m}48^{s}.07, δ = +44^{o}06′18′′.4 (equinox 2000.0), which is about 15′′ west and 56′′ south of the center of NGC 3938. Nothing is visible at this location on CCD frames taken by Rich on Feb. 2.18 and 27.16 (limiting mag ∼ 18.6) or on Palomar Digital Sky Survey images from 1997 (limiting red mag 20.0) and 1990 (limiting blue mag 19.5).

COMETS C/1997 B4 (SOHO) AND C/1997 V7 (SOHO) The initial observations are given below for two additional comets

found on archival SOHO website images (continuation to IAUC 8499). These objects do not belong to recognized groups, and their orbits (although uncertain) appear to be retrograde.

Comet  1997 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/1997 B4 Jan. 21.240 20^{h}02.^{m}8 −22^{o}36′ C3 RK 2005-F31
C/1997 V7 Nov. 2.529 14 26.8 −11 03 C3 RK 2005-F31
2005 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green